А K-100 N/m HA B 5

mA = 5kg
mB = 13kg
Uk= .2

1) Block B falls a distance of .6m. Find velocity of block A

2) Before it pauses for a moment what is Max Distance of B
Answer #1

T 139-T= 132 – 0 B a = 139 - a AT T-magsino-Umag aso - Ka = mala A =59 suisbunl 1oox = sq lo T-Sg83 0.2xSxgx5 - 100x T-39-0.8go 6 ( 3 (381 232 - Home ܘ ܘ ܐ 90.252 100x (0.62? x 0.6 Salon D) (8 18 - 3.0084 -1 VA 2 m/s le v=o 2 Xmax 100 go 23x max - I